# Predictive Assessment and Comparison of Bayesian Survival Models for Cancer Recurrence

### v1

@@ 5 Case Study: Gastrointestinal Stromal Tumour
However, we are not allowed to publish results from the real data set in this paper. Thus, we created a simulated data set that attempts to mimic the original data set and data sets that have been used in previous GIST studies (joensuu2012; joensuu2014). As the data are synthetic, the purpose of this section is not to focus on the results. Rather, the section aims to focus on the methodology that is presented in earlier sections of this paper and is now applied in practice.
The synthetic data set was generated using the following procedure. First, we removed data points with missing values from the original data set and fit the Bernoulli model to it. Then, we simulated covariates. We chose only covariates that were used by joensuu2012, although the original data set had other covariates as well. The selected covariates were Size, AgeAtSurg, MitHPF, GenderMale, Rupture, and Gastric. In addition, we included information on whether the patient received adjuvant treatment or not in a variable AdjTreatm. We created the synthetic data set consisting of these variables with the syn() function of the synthpop R package (gillian2016) using the default parameters. As can be seen in Figure 16, the distributions of the variables in the synthetic data set are very similar to the distributions of the variables in the original data set.
@@
After that, we created an artificial PatientID for each patient and started to produce rows with varying values on the time-dependent variables we created. For simplicity, we assumed that if the patient had received adjuvant treatment after surgery, it would always last three years. We created a data point for each year after surgery and predicted whether the recurrence occurred for that year using the Bernoulli model fitted to the original data. We introduced the same time-dependent variables as in joensuu2014: Time indicating the year after surgery, AdjOn indicating whether adjuvant treatment was in effect that year, and TimeSinceAdjStopped indicating how many years have passed since adjuvant treatment ended. We created predictions for consecutive years for each patient until cancer recurrence was predicted or 10 years had passed after surgery. In this case, we considered the event time to be censored. The prediction was stored as a binary variable Event. An example set of rows for one patient is shown in Table 1.
@@
Column Row 1 Row 2 Row 3 Row 4 Row 5
PatientID 60 60 60 60 60
Size 75 75 75 75 75
AgeAtSurg 64 64 64 64 64
MitHPF 13 13 13 13 13
GenderMale 0 0 0 0 0
Rupture 0 0 0 0 0
Gastric 1 1 1 1 1
Time 1 2 3 4 5
AdjOn 1 1 1 0 0
TimeSinceAdjStopped 0 0 0 1 2
Event 0 0 0 0 1
A detailed description of each variable is given below.
@@
We also generated a short form of the data set in order to be able to use AFT and PH models in addition to the Bernoulli model. The short form of the data set contains only one row for each patient and does not include time-dependent variables. Instead, adjuvant treatment receival is indicated with a single binary variable AdjTreatm, and the target variable is an integer EventTime, which indicates how many years after surgery the event occurs. We now also have a binary variable Censored, which indicates whether the patient is censored. The short form of the data for the patient in Table 1 is shown in Table 2.
@@
Column Row
PatientID 60
Size 75
AgeAtSurg 64
MitHPF 13
GenderMale 0
Rupture 0
Gastric 1
AdjTreatm 1
Censored 0
EventTime 5
@@ 5.1 Models
Before fitting any models, the continuous variables Size, AgeAtSurg, and MitHPF were scaled to have 0 mean and 0.5 standard deviation following the recommendation of gelman2008. This was done to make the regression coefficients of continuous variables comparable to the coefficients of binary variables. After scaling the variables, we fit four models: the Bernoulli model, the exponential model, the Cox proportional hazards model, and the Weibull accelerated failure time model. The Bernoulli model was fitted to the long data form, and the latter three models were fitted to the short data form. All priors were kept weakly information, since we did not have strong prior information available. Detailed description of the models is given below.
@@ 5.2 Posterior Predictive Checks
At the time of writing, the brms implementation of the Cox PH model is very limited. It lacks core functions, such as posterior_predict() and posterior_epred(). The implication of this is that posterior predictive checks among many other things cannot be performed for the Cox PH model. Since the implementation of these functions is outside the scope of this paper, we omit the posterior predictive check for the Cox PH model.
Since a large portion of the event times in the GIST data set are censored and the censoring process is not modelled, a suitable posterior predictive check for the parametric survival models (the exponential model and the Weibull AFT model) is the Kaplan-Meier overlay plot, where extrapolation is limited to 20%. The Kaplan-Meier overlay plot for the exponential model is shown in Figure 17. In the simulated data set, EventTime is a discrete variable, so the Kaplan-Meier survival curve for the observations is not very smooth. However, the figure shows that in all discrete time points $1,2,3,\dots,10$, the Kaplan-Meier estimate for the observations falls within the empirical CCDF estimates of the predictive draws. Thus, the predictive model check does not indicate a lack of fit for the exponential model.
@@
As the Bernoulli model has a binary outcome, the PAV-adjusted calibration plot is the recommended posterior predictive check for it. This is visualised in Figure 19. The calibration curve does not exceed the confidence bands, so there is no apparent lack of fit.
@@ 5.3 Predictive Model Comparison
Table 3: Result of model comparison with regular PSIS-LOO CV.
@@
Since the Weibull AFT model and the exponential model are parametric survival models, we can use their cumulative distribution functions to calculate the probability that an event happens before some fixed time point. For positive time points $t$, the CDF of the Weibull distribution is
@@
We predicted the probability that the event occurs at most 5 years after surgery, calculated log scores for the predicted probabilities, and performed a PSIS-LOO CV to compare the predictive performance of the Weibull AFT model and the exponential model. The result is presented in Table 4.
@@
The table shows that the exponential model outperforms the Weibull AFT model in this prediction task. It should be noted that, while the dichotomisation simplifies the prediction task, it loses quite a bit of information. As explained in Subsection 4.2, it is also possible to perform a comparison that loses less information while still staying in probabilities. This can be done by discretising time to some fixed length intervals and predicting the probability that an event occurs in each of the intervals. However, this approach is more complicated to implement.
Both the simple dichotomisation approach and the more complicated time discretisation approach allow the inclusion of the Bernoulli model in the comparison. However, cross-validation for the Bernoulli model requires leaving out all data points corresponding to one patient, which means that PSIS cannot be used to estimate the leave-one-out posterior. Thus, performing cross-validation for the Bernoulli model requires fitting the model separately for each patient, which is computationally very expensive.
That being said, in the GIST data, adjuvant treatment has a clear time-dependent effect on the probability of recurrence, which allows determining the superiority of the Bernoulli model without any computational comparison. This can be demonstrated visually. In Figure 21(b) and Figure 22(b), we visualise the predicted hazard rate from the exponential model and the Weibull AFT model, respectively. Similarly, Figure 23(b) illustrates the predicted recurrence probability from the Bernoulli model for each time interval. This can be thought of as a discrete time equivalent of the hazard rate.
@@
(a) Adjuvant treatment given.
(b) No adjuvant treatment given.
As can be seen in the figures, the Bernoulli model handles the time-dependent effect of adjuvant treatment on the recurrence probability in a way that is not possible in either the exponential model or the Weibull AFT model due to their model assumptions and inability to include variables with time-dependent effects. Since brms has a very limited implementation of the Cox PH model at the moment, we cannot visualise the predicted hazard from the Cox PH model. However, the proportional hazards assumption makes the Cox PH model incapable of capturing the time-dependent effect of the treatment visualised in Figure 23(b). Therefore, we can safely state that the Bernoulli model is superior to all other models presented in this paper when it comes to predicting the probability of recurrence of a gastrointestinal stromal tumour.
@@ 6.1 Predictive Model Checking
Scenario 1: Parametric survival models in the absence of censored event times If the data set does not contain data points with censored event times, many diagnostic plots, such as the intervals plot, the PIT-ECDF plot, and the Kaplan-Meier overlay plot, are applicable for assessing predictions of parametric survival models. If some observations are left-truncated, we recommend taking the truncation into account when plotting the survival curves in the Kaplan-Meier overlay plot.
Scenario 2: Parametric survival models in the presence of censored event times In case there are censored event times present and the censoring process is not modelled, the intervals plot and the PIT-ECDF plot become unusable, at least without the imputation of censored event times. The Kaplan-Meier overlay plot remains usable, but we recommend fine-tuning it by cutting the x-axis so that the plot extrapolates no more than 20% beyond the farthest observation. In addition, we recommend imputing censored event times and plotting a survival curve that includes them with a different colour than the survival curve for just the observed data points. The recommendation of taking left-truncation into account in the survival curves still applies.
Scenario 3: The Bernoulli model We recommend following the recommendation by sailynoja2025 and utilising the PAV-adjusted calibration plots. We extend their recommendation by suggesting to examine a zoomed plot in addition to the original one in case a large portion of predicted probabilities fall on a specific interval.
@@ 7.2 Future Directions
From the limitations of this paper we can naturally move on to potential future research directions. First of all, the research could be extended to survival models outside the brms framework. Secondly, in addition to the log scores, model comparison could focus on other scoring rules that directly address the issue with censored event times. These kinds of scoring rules include Survival-CRPS (avati2019) and strictly proper censoring-adjusted separable scoring rule (alberge2024), for example. Third, even though we implemented some of our findings in software tools, some remain unimplemented. Our recommendation to impute censored event times, when performing predictive model checks could be implemented to brms and bayesplot in the future. Software implementations could also be made to make it easier to compare models using our recommended method based on changing the prediction task so that all pointwise elpds are probabilities rather than densities. Fourth, as noticed while writing the paper, the Cox PH model does not have a predictive model check. This could be created and implemented. Finally, the methodology presented in this paper could be applied in the analysis of real data on gastrointestinal stromal tumours. This is something we plan to do in the future.
Acknowledgments

### v2

@@ 5 Case study: Gastrointestinal stromal tumour
However, we are not allowed to publish results from the real data set in this paper. Thus, we created a simulated data set that attempts to mimic the original data set and data sets that have been used in previous GIST studies (Joensuu et al., 2012, 2014). As the data are synthetic, the purpose of this section is not to focus on the results. Rather, the section aims to focus on the model checking methodology that is presented in earlier sections of this paper and is now applied in practice.
The synthetic data set was generated using the following procedure. First, we removed data points with missing values from the original data set and fit the Bernoulli model to it. Then, we simulated covariates. We chose only covariates that were used by Joensuu et al. (2012), although the original data set had other covariates as well. The selected covariates were Size, AgeAtSurg, MitHPF, GenderMale, Rupture, and Gastric. In addition, we included information on whether the patient received adjuvant treatment or not in a variable AdjTreatm. We created the synthetic data set consisting of these variables with the syn() function of the synthpop R package (Nowok et al., 2016) using the default parameters. As can be seen in Figure 16, the distributions of the variables in the synthetic data set are very similar to the distributions of the variables in the original data set.
@@
After that, we created an artificial PatientID for each patient and produced rows with varying values on the time-dependent variables. For simplicity, we assumed that if the patient had received adjuvant treatment after surgery, it would always last three years. We created a data point for each year after surgery and predicted whether the recurrence occurred for that year using the Bernoulli model fitted to the original data. We introduced the same time-dependent variables as in Joensuu et al. (2014): Time indicating the year after surgery, AdjOn indicating whether adjuvant treatment was in effect that year, and TimeSinceAdjStopped indicating how many years have passed since adjuvant treatment ended. We created predictions for consecutive years for each patient until cancer recurrence was predicted or 10 years had passed after surgery. In this case, we considered the event time to be censored. The prediction was stored as a binary variable Event. An example set of rows for one patient is shown in Table 1, and each variable is described in detail in Table 2.
PatientID Size AgeAtSurg MitHPF GenderMale Rupture Gastric Time AdjOn TimeSinceAdjStopped Event
60 75 64 13 0 0 1 1 1 0 0
60 75 64 13 0 0 1 2 1 0 0
60 75 64 13 0 0 1 3 1 0 0
60 75 64 13 0 0 1 4 0 1 0
60 75 64 13 0 0 1 5 0 2 1
@@
Variable Description
@@
Table 2: A detailed description of each variable.
We also generated a short form of the data set in order to be able to use AFT and PH models in addition to the Bernoulli model. The short form of the data set contains only one row for each patient and does not include time-dependent variables. Instead, adjuvant treatment receival is indicated with a single binary variable AdjTreatm, and the target variable is an integer EventTime, which indicates how many years after surgery the event occurs, as generated using the time-dependent variable model. We now also have a binary variable Censored, which indicates whether the patient is censored. The short form of the data for the patient in Table 1 is shown in Table 3.
PatientID Size AgeAtSurg MitHPF GenderMale Rupture Gastric AdjTreatm Censored EventTime
60 75 64 13 0 0 1 1 0 5
@@ 5.1 Models
Before fitting any models, the continuous variables Size, AgeAtSurg, and MitHPF were scaled to have 0 mean and 0.5 standard deviation following the recommendation of Gelman (2008). This was done to make the regression coefficients of continuous variables comparable to the coefficients of binary variables. After scaling the variables, we fit four models: the Bernoulli model, the exponential model, the Cox proportional hazards model, and the Weibull accelerated failure time model. The Bernoulli model was fitted to the long data form, and the latter three models were fitted to the short data form. We used weakly informative priors, since we did not have strong prior information available. Detailed description of the models is given below.
@@ 5.2 Posterior predictive checks
At the time of writing, the brms implementation of the Cox PH model is very limited. It lacks core functions, such as posterior_predict() and posterior_epred(). The implication of this is that posterior predictive checks among many other things cannot be performed for the Cox PH model. Since the implementation of these functions is outside the scope of this paper, we omit the posterior predictive checks for the Cox PH model.
Since a large portion of the event times in the GIST data set are censored and the censoring process is not modelled, a suitable posterior predictive check for the parametric survival models (the exponential model and the Weibull AFT model) is the Kaplan-Meier overlay plot, where the extrapolation is limited to 20% of the maximum event or censoring time. The Kaplan-Meier overlay plot for the exponential model is shown in Figure 17. In the simulated data set, EventTime is a discrete variable, so the Kaplan-Meier survival curve for the observations is not very smooth. However, the figure shows that in all discrete time points $1,2,3,\dots,10$, the Kaplan-Meier estimate for the observations falls within the empirical CCDF estimates of the predictive draws. Thus, the predictive model check does not indicate a lack of fit for the exponential model.
@@
As the Bernoulli model has a binary outcome, the PAV-adjusted calibration plot is the recommended posterior predictive check for it. This is visualised in Figure 19. The calibration curve does not exceed the consistency bands, so there is no apparent lack of fit.
@@ 5.3 Predictive Model Comparison
Table 4: Result of model comparison with regular PSIS-LOO CV.
@@
Since the Weibull AFT model and the exponential model are parametric survival models, we can use their cumulative distribution functions to compute the probability that an event happens before some fixed time point. For positive time points $t$, the CDF of the Weibull distribution is
@@
We predicted the probability that the event occurs at most 5 years after surgery, computed log scores for the predicted probabilities, and performed a PSIS-LOO CV to compare the predictive performance of the Weibull AFT model and the exponential model. The result is presented in Table 5.
@@
The table shows that the exponential model outperforms the Weibull AFT model in this prediction task. It should be noted that, while the dichotomisation simplifies the prediction task, it loses quite a bit of information. As explained in Subsection 4.2, it is also possible to perform a comparison that loses less information while still comparing predictive probabilities. This can be done by discretising time to some fixed length intervals and predicting the probability that an event occurs in each of the intervals. However, this approach is more complicated to implement.
Both the simple dichotomisation approach and the more complicated time discretisation approach allow the inclusion of the Bernoulli model in the comparison. Cross-validation for the Bernoulli model requires leaving out all data points corresponding to one patient and PSIS may fail as the posterior of a flexible model can change too much for importance weighting based approach to work reliably. Performing cross-validation for the Bernoulli model may require fitting the model separately for each patient, which can be computationally expensive.
That being said, in the GIST data, adjuvant treatment has a clear time-dependent effect on the probability of recurrence, which allows determining the superiority of the Bernoulli model without any computational comparison. This can be demonstrated visually. In Figure 21 and Figure 22, we visualise the predicted hazard rate from the exponential model and the Weibull AFT model, respectively. Similarly, Figure 23 illustrates the predicted recurrence probability from the Bernoulli model for each time interval. This can be thought of as a discrete time equivalent of the hazard rate.
@@
(a) Adjuvant treatment given.
(b) No adjuvant treatment given.
@@
As can be seen in the figures, the Bernoulli model handles the time-dependent effect of adjuvant treatment on the recurrence probability in a way that is not possible in either the exponential model or the Weibull AFT model due to their model assumptions and inability to include variables with time-dependent effects. Since brms has a very limited implementation of the Cox PH model at the moment, we cannot visualise the predicted hazard from the Cox PH model. However, the proportional hazards assumption makes the Cox PH model incapable of capturing the time-dependent effect of the treatment visualised in Figure 23. Therefore, we can safely state that the Bernoulli model is superior to all other models presented in this paper when it comes to predicting the probability of recurrence of a gastrointestinal stromal tumour.
@@ 6.1 Predictive model checking
Scenario 1: Parametric survival models in the absence of censored event times If the data set does not contain censored event times, many diagnostic plots, such as the intervals plot, the PIT-ECDF plot, and the Kaplan-Meier overlay plot, are applicable for assessing predictions of parametric survival models. If some observations are left-truncated, we recommend taking the truncation into account when plotting the survival curves in the Kaplan-Meier overlay plot.
Scenario 2: Parametric survival models in the presence of censored event times In case there are censored event times and the censoring process is not modelled, the intervals plot and the PIT-ECDF plot become unusable, at least without the imputation of censored event times. The Kaplan-Meier overlay plot remains usable, but we recommend fine-tuning it by cutting the x-axis so that the plot extrapolates no more than 20% beyond the furthest observation. In addition, we recommend imputing censored event times and plotting a survival curve that includes them with a different colour than the survival curve for just the observed data points. The recommendation of taking left-truncation into account in the survival curves still applies.
Scenario 3: The Bernoulli model We recommend following the recommendation by Säilynoja et al. (2025) and utilising the PAV-adjusted calibration plots. We extend their recommendation by suggesting to examine a zoomed plot in addition to the original one in case a large portion of predicted probabilities fall on a specific interval.
@@ 7.2 Future Directions
From the limitations of this paper we can naturally move on to potential future research directions. First of all, the research could be extended to survival models outside the brms framework. Secondly, in addition to the log scores, model comparison could focus on other scoring rules that directly address the issue with censored event times. These kinds of scoring rules include Survival-CRPS (Avati et al., 2019) and strictly proper censoring-adjusted separable scoring rule (Alberge et al., 2024). Third, even though we implemented some of our findings in software tools, some remain unimplemented. Our recommendation to impute censored event times, when performing predictive model checks could be implemented to brms and bayesplot in the future. Software implementations could also be made to make it easier to compare models using our recommended method based on changing the prediction task so that all pointwise elpds are probabilities rather than densities. Fourth, as noticed while writing the paper, the Cox PH model does not have a predictive model check. This could be created and implemented. Finally, the methodology presented in this paper could be applied in the analysis of real data on gastrointestinal stromal tumours.
Acknowledgements